### v1

# HGMF: A Hierarchical Gaussian Mixture Framework for Scalable Tool Invocation within the Model Context Protocol

Abstract
Invoking external tools enables Large Language Models (LLMs) to perform complex, real-world tasks, yet selecting the correct tool from large, hierarchically-structured libraries remains a significant challenge. The limited context windows of LLMs and noise from irrelevant options often lead to low selection accuracy and high computational costs. To address this, we propose the Hierarchical Gaussian Mixture Framework (HGMF), a probabilistic pruning method for scalable tool invocation. HGMF first maps the user query and all tool descriptions into a unified semantic space. The framework then operates in two stages: it clusters servers using a Gaussian Mixture Model (GMM) and filters them based on the query’s likelihood. Subsequently, it applies the same GMM-based clustering and filtering to the tools associated with the selected servers. This hierarchical process produces a compact, high-relevance candidate set, simplifying the final selection task for the LLM. Experiments on a public dataset show that HGMF significantly improves tool selection accuracy while reducing inference latency, confirming the framework’s scalability and effectiveness for large-scale tool libraries.
Index Terms—  Large Language Models, Tool Invocation, Hierarchical Clustering, Gaussian Mixture Model, Semantic Search
1 Introduction
The capacity of Large Language Models (LLMs) can be substantially extended by enabling them to interact with external tools, a process often referred to as tool invocation. This allows LLMs to perform complex, real-world tasks by accessing APIs, querying databases, or executing code. A fundamental prerequisite for successful tool invocation is the model’s ability to select the appropriate tool from a given library in response to a user’s request.
While effective for small, curated sets of tools, this selection process becomes a significant bottleneck when the tool library is large and hierarchically organized. Modern systems may contain thousands of tools nested within various servers or services. Presenting this entire collection to an LLM is often infeasible due to three primary constraints. First, the limited context window of most LLMs prohibits the inclusion of extensive tool descriptions. Second, even if the context window were sufficient, the presence of a vast number of irrelevant tools introduces significant semantic noise, which degrades the model’s reasoning capabilities and leads to a sharp decline in selection accuracy. Third, processing such large contexts results in prohibitive computational costs and high inference latency, rendering the system impractical for real-time applications.
Existing approaches to mitigate this issue typically rely on a preliminary filtering stage before the LLM’s final selection. Methods such as keyword matching or standard vector similarity search can reduce the candidate pool, but they often fail to capture the nuanced semantics of a user’s query. Furthermore, these methods typically treat the tool library as a flat list, thereby ignoring the valuable structural information inherent in server-tool hierarchies. This oversight can lead to the filtering out of relevant tools or the inclusion of contextually inappropriate ones.
To address these limitations, we propose the Hierarchical Gaussian Mixture Framework (HGMF), a novel, two-stage probabilistic pruning framework designed for scalable and accurate tool selection. HGMF operates by first embedding all textual descriptions into a shared semantic space. It then applies a hierarchical filtering process:
  • At the server level, it models the distribution of servers using a Gaussian Mixture Model (GMM) and identifies a small set of relevant server clusters based on the probabilistic likelihood of the user’s query.
  • At the tool level, it repeats this GMM-based clustering and filtering process exclusively for the tools associated with the pre-selected servers.
This hierarchical, probabilistic approach produces a compact and highly relevant candidate set, which is then passed to the LLM for a simplified and more accurate final selection. The main contributions of this work are as follows:
  • We propose HGMF, a novel hierarchical framework that effectively addresses the scalability challenge in LLM-based tool selection by structuring the filtering process according to the natural server-tool hierarchy.
  • We introduce the use of Gaussian Mixture Models for probabilistic, soft-clustering-based filtering, which more accurately captures the semantic relationship between the user query and clusters of tools compared to deterministic methods.
  • We demonstrate through empirical evaluation on a public dataset that HGMF significantly improves tool selection accuracy while concurrently reducing inference latency, establishing its effectiveness for large-scale applications.
2 Related Works
2.1 Tool-Augmented LLMs
Early large language models were primarily based on pre-training and fine-tuning to process natural language tasks, lacking the ability to interact with external entities, which limited their performance on tasks[1][2]. However, tool-enhanced large language models improve the model’s language reasoning performance in complex tasks through external tools and databases. Different aspects of the problem have been addressed at different stages of development. RAG dynamically retrieves external data, allowing the model to obtain information from external documents. By combining retrieval and generation, it enhances the model’s contextual relevance and accuracy[3]. As large language models continue to grow in scale, WebGPT [4] enhances the real-time information processing capabilities of large models through interaction with web browsers.Toolformer [5] uses self-supervised learning methods to enable models to learn how to call external tools.While API calls to external tools can be used to extend LLM functionality, allowing specific tools to be selected based on task requirements, tool selection and interaction methods require manual design. LangChain [6]provides a modular framework for automatic tool selection, supporting the integration of large language models (LLMs) with external tools and data.Gorilla [7]proposed that LLM, which is optimized for API calls, can automatically select tools based on user queries. To address the issue of incompatible protocols across different platforms, standardized interfaces such as the Model Context Protocol (MCP) [8]have been proposed to unify the interaction between LLMs and external tools and databases. The open-source tool ecosystem and enterprise-level APIs have further driven the adoption of tool-enhanced LLMs.
2.2 Sampling Methods Combined With Large Language Models
With the development of large language models (LLM), combining sampling methods with model inference processes has emerged as a new research direction aimed at improving the accuracy and efficiency of model outputs. Temperature sampling controls the randomness of outputs by adjusting the smoothness of the probability distribution; higher temperature values increase the selection probability of low-probability words [9]. Top-K sampling selects only from the top K words with the highest probabilities, ensuring output accuracy but potentially compromising diversity. Subsequent research introduced Top-P sampling [10], which dynamically selects the smallest set of words with cumulative probabilities reaching P, balancing diversity and rationality. This has become the mainstream sampling strategy in LLM generation tasks. To improve problem-solving performance, prompt-based sampling optimization methods have been proposed. High-quality samples used as prompts can significantly enhance the model’s ability to perform specific tasks. The relevance of examples to the target task has a significant impact on sampling effectiveness. Therefore, some studies combine semantic similarity retrieval to sample the most relevant examples from an example repository and construct prompts, further enhancing the LLM’s reasoning capabilities [11]. Existing sampling methods still have limitations when handling complex scenarios: in large sample sizes, the semantic boundaries of samples are unclear, making it difficult to match users’ true intentions; descriptions of unrelated tools may contain similar vocabulary; sample quality is inconsistent; and large sample sizes can cause truly relevant tools to be overlooked. To address these issues, the sampling method proposed in this paper aims to enhance the adaptability and effectiveness of sampling in large sample sizes through user-semantics-driven intelligent sampling.
2.3 Model Context Protocol
Currently, large language models (LLMs) struggle to fully realize their problem-solving potential due to data limitations. To address this, we propose the Model Context Protocol (MCP), designed to enable efficient integration between LLMs and external databases and tools. MCP provides a standardized interface for connecting LLMs to diverse data sources and tools, addressing the pain point of having to customize different protocols for different platforms. By adopting a unified approach, MCP significantly enhances system interoperability and the efficiency of LLM invocations [8]. However, the current common method of obtaining tools through MCP relies on matching random samples with large language model outputs. This method has obvious shortcomings in practical applications: randomly sampled tools are often irrelevant to user queries and have low hit rates. This is especially true in large sample scenarios with a large number of tools, where the accuracy of matching declines significantly as the number of tools increases [12]. To address this issue, we propose a two-stage retrieval-based sampling method based on semantic vectors and GMM clustering to ensure that the selected tools are more closely aligned with the actual needs of user queries. Experiments show that this method significantly improves the accuracy of tool matching and effectively overcomes the limitations of traditional random sampling methods.
3 Methodology
In this section, we present our proposed framework, Hierarchical Gaussian Mixture Framework (HGMF). The core objective of HGMF is to efficiently prune a vast and hierarchically structured toolset to a small, highly relevant candidate pool. This refined pool is then presented to a Large Language Model (LLM) for final, precise tool selection. Our framework effectively mitigates the challenges of context window limitations and noise interference in LLM-based tool invocation. The overall architecture of HGMF is depicted in Fig. 1 and comprises three main stages: (1) Semantic Embedding and Preprocessing, (2) Hierarchical GMM-based Candidate Pruning, and (3) LLM-based Reranking and Selection.
3.1 Semantic Embedding and Preprocessing
The initial stage aims to represent all textual information—the user query, server descriptions, and tool descriptions—within a unified high-dimensional semantic space. Let the set of all servers be $\mathcal{S}=\{s_{1},s_{2},\dots,s_{M}\}$ and the set of all tools be $\mathcal{T}=\{t_{1},t_{2},\dots,t_{N}\}$, where each tool $t_{j}$ is associated with a specific server $s_{i}$. Given a user query $q$, we first employ a pre-trained sentence transformer model, specifically all-MiniLM-L6-v2, to encode all textual descriptions into $d$-dimensional embedding vectors (where $d=384$). This process yields a query vector $\mathbf{v}_{q}$, a set of server vectors $\{\mathbf{v}_{s_{i}}\}_{i=1}^{M}$, and a set of tool vectors $\{\mathbf{v}_{t_{j}}\}_{j=1}^{N}$. To ensure that the subsequent similarity and probability calculations are not biased by vector magnitudes, we apply L2 normalization to all generated embeddings. For any vector $\mathbf{v}$, its normalized counterpart $\hat{\mathbf{v}}$ is computed as:
$\hat{\mathbf{v}}=\frac{\mathbf{v}}{\|\mathbf{v}\|_{2}}$ (1)
This normalization projects all vectors onto the surface of a unit hypersphere, making cosine similarity equivalent to the dot product and stabilizing the clustering process.
Fig. 1: Overall Pipeline. The process involves embedding the user query, servers, and tools into a unified semantic space. HGMF then performs a two-level pruning strategy using clustering at the server level to obtain a relevant server list, followed by clustering at the tool level within the selected servers to yield a refined list of candidate toolsets.
3.2 Hierarchical GMM-based Candidate Pruning
This stage is the cornerstone of our method, designed to drastically reduce the search space by exploiting the inherent server-tool hierarchy. We employ a two-level pruning strategy using Gaussian Mixture Models (GMMs).
Server-Level Pruning.
First, we model the distribution of all server embeddings $\{\hat{\mathbf{v}}_{s_{i}}\}_{i=1}^{M}$ using a GMM. A GMM assumes that the data is generated from a mixture of a finite number of Gaussian distributions with unknown parameters. We fit a GMM with $K_{s}$ components to the server vectors. A common heuristic, which we adopt, is to set $K_{s}=\lceil\sqrt{M}\rceil$ to balance model complexity and expressiveness. After fitting the model, we obtain $K_{s}$ Gaussian components, each characterized by a mean (cluster centroid) $\bm{\mu}_{k}$ and a covariance matrix $\bm{\Sigma}_{k}$. We then evaluate the relevance of each server cluster to the user query by calculating the likelihood of the query vector $\hat{\mathbf{v}}_{q}$ under each Gaussian component. The likelihood score for the $k$-th cluster is given by its probability density function:
$\mathcal{L}(k|\hat{\mathbf{v}}_{q})=\mathcal{N}(\hat{\mathbf{v}}_{q}|\bm{\mu}_{k},\bm{\Sigma}_{k})$ (2)
We rank all server clusters based on these likelihood scores in descending order and select the top-$N_{s}$ clusters. All servers belonging to these selected clusters form our pruned server set, $\mathcal{S}^{\prime}\subset\mathcal{S}$.
Tool-Level Pruning.
Next, for each server $s_{i}\in\mathcal{S}^{\prime}$, we consider its associated set of tool embeddings:
$\mathcal{T}_{i}=\left\{\hat{\mathbf{v}}_{t_{j}}\,\middle|\,t_{j}\text{ is associated with }s_{i}\right\}$
We independently fit a new GMM with $K_{t}=\lceil\sqrt{|\mathcal{T}_{i}|}\rceil$ components to each tool set $\mathcal{T}_{i}$. Similar to the server-level pruning, we compute the likelihood of the query vector $\hat{\mathbf{v}}_{q}$ for each tool cluster within each selected server. For each server $s_{i}$, we rank its tool clusters and select the top-$N_{t}$ clusters. The tools belonging to these top-ranked clusters form the final pruned candidate tool set, $\mathcal{T}^{\prime}\subset\mathcal{T}$. This hierarchical process ensures that only tools associated with relevant servers are considered, leading to a highly refined and contextually coherent candidate pool.
Fig. 2: Performance comparison of our proposed method (HGMF) against five baselines across eight different Large Language Models (LLMs). Each subplot corresponds to a specific LLM, showing accuracy (Y-axis) as a function of sample size (X-axis, log scale). The results indicate that HGMF achieves state-of-the-art performance under most models and shot counts, with a particularly significant advantage at larger sample sizes.
3.3 LLM-based Reranking and Final Selection
After pruning, we have a small, structured candidate set $\{(s_{i},t_{j})|s_{i}\in\mathcal{S}^{\prime},t_{j}\in\mathcal{T}^{\prime}\}$. This set is now small enough to be efficiently processed by an LLM.
LLM-driven Description Generation.
We construct a structured prompt that presents the pruned candidate set to the LLM and instructs it to act as a helpful assistant. The LLM’s task is to analyze the user query $q$ and generate a natural language description of the ideal server and the ideal tool required to fulfill the request. This step leverages the LLM’s superior reasoning and language understanding capabilities to synthesize a ”perfect” target description based on the provided high-quality candidates.
Final Matching and Scoring.
Let the LLM-generated descriptions for the server and tool be $d_{s}^{*}$ and $d_{t}^{*}$, respectively. We encode these descriptions using the same sentence transformer model to obtain their embeddings, $\mathbf{v}_{s}^{*}$ and $\mathbf{v}_{t}^{*}$. We then perform the final matching by computing the cosine similarity between the LLM-generated embeddings and the embeddings of all candidates in the pruned set. The server score for a candidate server $s_{i}\in\mathcal{S}^{\prime}$ is:
$\text{score}(s_{i})=\cos(\mathbf{v}_{s}^{*},\hat{\mathbf{v}}_{s_{i}})$ (3)
Similarly, the tool similarity for a candidate tool $t_{j}\in\mathcal{T}^{\prime}$ is:
$\text{sim}(t_{j})=\cos(\mathbf{v}_{t}^{*},\hat{\mathbf{v}}_{t_{j}})$ (4)
To determine the final ranking, we compute a combined score for each server-tool pair $(s_{i},t_{j})$ that accounts for both server-level and tool-level matching quality. The final score is calculated as:
$\begin{split}\text{FinalScore}(s_{i},t_{j})={}&(\text{score}(s_{i})\times\text{sim}(t_{j}))\\ &\times\max(\text{score}(s_{i}),\text{sim}(t_{j}))\end{split}$ (5)
This formula gives higher weight to pairs where both the server and tool are a strong match, and it particularly rewards candidates where at least one component has an exceptionally high similarity score. The server-tool pair with the highest final score is selected as the output of our HGMF framework.
4 Experiments
4.1 Experiment Configuration
Dataset
The dataset employed in this study is MCP-tools, introduced by Fei et al. [13]. It comprises 2,797 tools collected from 308 servers within the official MCP repository. For our semantic matching task, we preprocessed all textual content by generating vector embeddings using the all-MiniLM-L6-v2 sentence-transformer model [14]. The original data, along with these pre-computed embeddings, were consolidated and stored in a single JSON file.
Table 1: Average accuracy (%) of all methods across different LLMs. The best result in each column is in bold. Full results are in the Appendix.
Method Mistral (7b) Gemma (7b) Llama3 (8b) Qwen (14b) Solar (10.7b) Openhermes (7b) Phi3 (3.8b) Phi4 (14b) Average
MCP-zero 80.95 79.52 82.74 86.26 80.36 75.79 59.48 79.70 78.10
HGMF (Ours) 91.32 79.19 84.83 87.21 87.99 82.57 62.58 82.99 82.34
Tokenization-Based 58.65 38.43 67.50 79.88 69.61 47.48 40.92 80.18 60.32
K-means 82.75 62.05 83.52 87.51 75.85 74.15 54.80 84.95 75.70
Cluster-Weighted 73.21 56.04 74.73 84.25 78.28 71.87 55.74 82.93 72.13
Density-Based 78.35 56.77 83.15 78.64 75.56 60.87 55.86 79.60 71.10
Implementation Details
Our implementation uses Python 3.10 and PyTorch 2.7.1 (CUDA 12.6), executed on NVIDIA RTX 4090 GPUs. Key hyperparameters are: $k=4$ for top-k retrieval, a maximum of 10 tools indexed per server, and random sampling for sample sizes under 10 to mitigate noise. We evaluate on 10 exponentially scaled sample sizes ranging from 1 to 2797.
Evaluation Metrics
We evaluate model performance using Exact Match Accuracy. A prediction is considered correct only if both the retrieved server and tool perfectly match the ground truth. The metric is calculated as:
$\text{Accuracy}=\frac{\text{\# Correct Predictions}}{\text{\# Total Test Cases}}$
We report accuracy for the entire test set (Overall) and for each individual sample size tier ($s$).
Baselines.
We compare our method against the following baselines:
  • MCP-zero [13]: The current state-of-the-art, which uses random sampling without replacement.
  • Tokenization-Based: A naive approach that samples based on tokenizing the input.
  • K-means: A standard clustering approach where we sample from the $k$ clusters closest to the query embedding.
  • Cluster-Weighted: An extension of K-means that assigns sampling weights to tools within selected clusters based on their cosine similarity to the query.
  • Density-Based Clustering: A method that groups servers and tools based on their intrinsic embedding similarities without a predefined number of clusters.
Fig. 3: Accuracy gain of the HGMF method over the MCP-zero baseline. The results show that HGMF’s performance advantage steadily increases with the sample size, demonstrating its superior scalability.
4.2 Result Analysis
Table 1 and Fig. 2 present a comprehensive comparison of our proposed method, HGMF, against five baseline methods across eight different LLMs and four sample sizes. The results demonstrate that HGMF consistently achieves superior accuracy, particularly in high-shot settings. Fig. 3 shows the accuracy improvement of our proposed HGMF method over the MCP-zero baseline across eight LLMs and four sample sizes. Positive values indicate performance gains, while negative values indicate underperformance. Key observations are summarized as follows:
Overall Performance.
HGMF attains the highest average accuracy (82.34%) across all models, outperforming the strong baseline MCP-zero (78.10%) by over 4 percentage points. It achieves the best results on 6 out of 8 models, including Mistral:7b, Llama3:8b, Qwen:14b, Solar:10.7b, Openhermes:7b, and Phi3:3.8b. This confirms the effectiveness of our hierarchical clustering and pruning strategy in selecting high-quality tool candidates.
Comparison with Clustering-Based Baselines:
While K-means and Density-Based Clustering exhibit moderate improvements over Tokenization-Based sampling, they fall short of HGMF. Notably, Cluster-Weighted performs inconsistently, likely due to its reliance on local similarity scores without hierarchical context. In contrast, HGMF integrates both server-level and tool-level clustering, enabling more robust and context-aware selection.
Scalability with Sample Size:
HGMF’s performance gain increases consistently with more samples—from 41-shot (light gray) to 2797-shot (dark blue)—demonstrating strong scalability and its ability to extract high-quality candidates from larger tool pools.
Low-Shot Limitations:
In 41-shot settings, HGMF may underperform due to insufficient context for effective clustering. Notably, it lags behind MCP-zero on models like Gemma:7b and phi4:14b, with up to 30% accuracy drop.
High-Shot Superiority:
At larger sample sizes ($\geq$401-shot), HGMF consistently outperforms the baseline. At 2797-shot, it delivers significant gains—over 40 percentage points on Openhermes:7b and 30+ points on phi4:14b, highlighting its strength in distilling relevant tools from noisy candidates.
Model Sensitivity:
Response to HGMF varies by model. Gemma:7b shows large fluctuation, while models like Mistral:7b and Qwen:14b exhibit stable improvements across all settings.
5 Conclusion
To address the challenges of context limitations and noise interference when Large Language Models select tools from large, hierarchical libraries, this paper introduces the Hierarchical Gaussian Mixture Framework (HGMF). The framework leverages the library’s inherent ”server-tool” structure, performing efficient probabilistic pruning on the vast toolset through two-stage Gaussian Mixture Model (GMM) clustering to produce a compact, high-relevance candidate set for the LLM.
Experiments confirm HGMF’s effectiveness and scalability. It achieves higher accuracy than baseline methods like random sampling across multiple LLMs. Critically, its performance advantage grows with the size of the tool library, demonstrating that the framework can effectively handle the challenges of large-scale, real-world toolsets. Through this intelligent filtering, HGMF not only improves selection accuracy but also reduces the LLM’s inference burden.
The core contribution of this work is an efficient and viable paradigm for large-scale tool invocation. By integrating hierarchical information into the pruning process, HGMF successfully decomposes a complex selection problem into more manageable sub-tasks. Although its performance is limited with very small toolsets, this work points toward future research, such as exploring adaptive clustering parameters and extending the framework to more complex, graph-structured tool libraries.

### v2

@@
 HGMF: A Hierarchical Gaussian Mixture Framework for Scalable Tool Invocation within the Model Context Protocol
-Abstract
+Abstract.
 Invoking external tools enables Large Language Models (LLMs) to perform complex, real-world tasks, yet selecting the correct tool from large, hierarchically-structured libraries remains a significant challenge.
@@
 Experiments on a public dataset show that HGMF significantly improves tool selection accuracy while reducing inference latency, confirming the framework’s scalability and effectiveness for large-scale tool libraries.
-Index Terms—  Large Language Models, Tool Invocation, Hierarchical Clustering, Gaussian Mixture Model, Semantic Search
-1 Introduction
-The capacity of Large Language Models (LLMs) can be substantially extended by enabling them to interact with external tools, a process often referred to as tool invocation.
-This allows LLMs to perform complex, real-world tasks by accessing APIs, querying databases, or executing code.
-A fundamental prerequisite for successful tool invocation is the model’s ability to select the appropriate tool from a given library in response to a user’s request.
-While effective for small, curated sets of tools, this selection process becomes a significant bottleneck when the tool library is large and hierarchically organized.
-Modern systems may contain thousands of tools nested within various servers or services.
-Presenting this entire collection to an LLM is often infeasible due to three primary constraints.
-First, the limited context window of most LLMs prohibits the inclusion of extensive tool descriptions.
-Second, even if the context window were sufficient, the presence of a vast number of irrelevant tools introduces significant semantic noise, which degrades the model’s reasoning capabilities and leads to a sharp decline in selection accuracy.
-Third, processing such large contexts results in prohibitive computational costs and high inference latency, rendering the system impractical for real-time applications.
-Existing approaches to mitigate this issue typically rely on a preliminary filtering stage before the LLM’s final selection.
-Methods such as keyword matching or standard vector similarity search can reduce the candidate pool, but they often fail to capture the nuanced semantics of a user’s query.
-Furthermore, these methods typically treat the tool library as a flat list, thereby ignoring the valuable structural information inherent in server-tool hierarchies.
-This oversight can lead to the filtering out of relevant tools or the inclusion of contextually inappropriate ones.
-To address these limitations, we propose the Hierarchical Gaussian Mixture Framework (HGMF), a novel, two-stage probabilistic pruning framework designed for scalable and accurate tool selection.
-HGMF operates by first embedding all textual descriptions into a shared semantic space.
-It then applies a hierarchical filtering process:
-  • At the server level, it models the distribution of servers using a Gaussian Mixture Model (GMM) and identifies a small set of relevant server clusters based on the probabilistic likelihood of the user’s query.
-  • At the tool level, it repeats this GMM-based clustering and filtering process exclusively for the tools associated with the pre-selected servers.
-This hierarchical, probabilistic approach produces a compact and highly relevant candidate set, which is then passed to the LLM for a simplified and more accurate final selection.
-The main contributions of this work are as follows:
-  • We propose HGMF, a novel hierarchical framework that effectively addresses the scalability challenge in LLM-based tool selection by structuring the filtering process according to the natural server-tool hierarchy.
-  • We introduce the use of Gaussian Mixture Models for probabilistic, soft-clustering-based filtering, which more accurately captures the semantic relationship between the user query and clusters of tools compared to deterministic methods.
-  • We demonstrate through empirical evaluation on a public dataset that HGMF significantly improves tool selection accuracy while concurrently reducing inference latency, establishing its effectiveness for large-scale applications.
-2 Related Works
-2.1 Tool-Augmented LLMs
-Early large language models were primarily based on pre-training and fine-tuning to process natural language tasks, lacking the ability to interact with external entities, which limited their performance on tasks[1][2].
-However, tool-enhanced large language models improve the model’s language reasoning performance in complex tasks through external tools and databases.
-Different aspects of the problem have been addressed at different stages of development.
-RAG dynamically retrieves external data, allowing the model to obtain information from external documents.
-By combining retrieval and generation, it enhances the model’s contextual relevance and accuracy[3].
-As large language models continue to grow in scale, WebGPT [4] enhances the real-time information processing capabilities of large models through interaction with web browsers.Toolformer [5] uses self-supervised learning methods to enable models to learn how to call external tools.While API calls to external tools can be used to extend LLM functionality, allowing specific tools to be selected based on task requirements, tool selection and interaction methods require manual design.
-LangChain [6]provides a modular framework for automatic tool selection, supporting the integration of large language models (LLMs) with external tools and data.Gorilla [7]proposed that LLM, which is optimized for API calls, can automatically select tools based on user queries.
-To address the issue of incompatible protocols across different platforms, standardized interfaces such as the Model Context Protocol (MCP) [8]have been proposed to unify the interaction between LLMs and external tools and databases.
-The open-source tool ecosystem and enterprise-level APIs have further driven the adoption of tool-enhanced LLMs.
-2.2 Sampling Methods Combined With Large Language Models
-With the development of large language models (LLM), combining sampling methods with model inference processes has emerged as a new research direction aimed at improving the accuracy and efficiency of model outputs.
-Temperature sampling controls the randomness of outputs by adjusting the smoothness of the probability distribution; higher temperature values increase the selection probability of low-probability words [9].
-Top-K sampling selects only from the top K words with the highest probabilities, ensuring output accuracy but potentially compromising diversity.
-Subsequent research introduced Top-P sampling [10], which dynamically selects the smallest set of words with cumulative probabilities reaching P, balancing diversity and rationality.
-This has become the mainstream sampling strategy in LLM generation tasks.
-To improve problem-solving performance, prompt-based sampling optimization methods have been proposed.
-High-quality samples used as prompts can significantly enhance the model’s ability to perform specific tasks.
-The relevance of examples to the target task has a significant impact on sampling effectiveness.
-Therefore, some studies combine semantic similarity retrieval to sample the most relevant examples from an example repository and construct prompts, further enhancing the LLM’s reasoning capabilities [11].
-Existing sampling methods still have limitations when handling complex scenarios: in large sample sizes, the semantic boundaries of samples are unclear, making it difficult to match users’ true intentions; descriptions of unrelated tools may contain similar vocabulary; sample quality is inconsistent; and large sample sizes can cause truly relevant tools to be overlooked.
-To address these issues, the sampling method proposed in this paper aims to enhance the adaptability and effectiveness of sampling in large sample sizes through user-semantics-driven intelligent sampling.
-2.3 Model Context Protocol
-Currently, large language models (LLMs) struggle to fully realize their problem-solving potential due to data limitations.
-To address this, we propose the Model Context Protocol (MCP), designed to enable efficient integration between LLMs and external databases and tools.
-MCP provides a standardized interface for connecting LLMs to diverse data sources and tools, addressing the pain point of having to customize different protocols for different platforms.
-By adopting a unified approach, MCP significantly enhances system interoperability and the efficiency of LLM invocations [8].
-However, the current common method of obtaining tools through MCP relies on matching random samples with large language model outputs.
-This method has obvious shortcomings in practical applications: randomly sampled tools are often irrelevant to user queries and have low hit rates.
-This is especially true in large sample scenarios with a large number of tools, where the accuracy of matching declines significantly as the number of tools increases [12].
-To address this issue, we propose a two-stage retrieval-based sampling method based on semantic vectors and GMM clustering to ensure that the selected tools are more closely aligned with the actual needs of user queries.
-Experiments show that this method significantly improves the accuracy of tool matching and effectively overcomes the limitations of traditional random sampling methods.
-3 Methodology
-In this section, we present our proposed framework, Hierarchical Gaussian Mixture Framework (HGMF).
-The core objective of HGMF is to efficiently prune a vast and hierarchically structured toolset to a small, highly relevant candidate pool.
-This refined pool is then presented to a Large Language Model (LLM) for final, precise tool selection.
-Our framework effectively mitigates the challenges of context window limitations and noise interference in LLM-based tool invocation.
-The overall architecture of HGMF is depicted in Fig.
-1 and comprises three main stages: (1) Semantic Embedding and Preprocessing, (2) Hierarchical GMM-based Candidate Pruning, and (3) LLM-based Reranking and Selection.
-3.1 Semantic Embedding and Preprocessing
+Abstract.
+Invoking external tools empowers large language models (LLMs) to execute complex, real-world tasks, yet selecting the appropriate tool from expansive, hierarchically structured libraries poses a formidable challenge.
+LLM context limitations and semantic noise from irrelevant options frequently yield suboptimal accuracy and elevated computational overhead.
+To mitigate these issues, we introduce the Hierarchical Gaussian Mixture Framework (HGMF), a probabilistic pruning paradigm for scalable tool invocation under the Model Context Protocol.
+HGMF begins by embedding user queries, servers, and tools into a unified semantic space via sentence transformers with L2 normalization.
+It then employs a two-stage GMM-based filtering: server-level clustering identifies query-relevant clusters via likelihood maximization, pruning to a subset; tool-level GMMs, applied within selected servers, refine candidates.
+To enhance robustness in low-sample regimes—where cluster boundaries blur and distributions destabilize—we integrate inter-class regularization (promoting separation between cluster centers) and intra-class regularization (ensuring compactness within clusters), balanced with covariance stabilization.
+The resultant compact, high-fidelity candidate pool facilitates LLM-driven reranking and precise final selection.
+Evaluations on the MCP-tools dataset across eight LLMs demonstrate HGMF’s superiority, outperforming over baselines like MCP-zero.
+Our code and data is available at xxx.
+Large Language Models, Tool Invocation, Hierarchical Clustering, Gaussian Mixture Model, Semantic Search
+$c$s: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
+1. Introduction
+While Large Language Models (LLMs) demonstrate exceptional performance in language tasks, their real-world utility remains constrained without integration with external knowledge resources.
+Concurrently, the deployment of these models faces significant security and legal hurdles.
+On the security front, LLMs are susceptible to sophisticated adversarial attacks and latent manipulations that compromise their reliability (Xing et al., 2025a; Li et al., 2025; Xing et al., 2025b).
+Meanwhile, the necessity for intellectual property safeguards has spurred research into robust copyright protection and tracing mechanisms to mitigate unauthorized use and data leakage (Xu et al., 2025b; Yue et al., 2025; Xu et al., 2025d, c; Zhang et al., 2025).
+To extend their functional boundaries, integrating external tools such as APIs and databases—a paradigm known as tool invocation—is essential for executing complex, multi-step operations (Schick et al., 2023).
+However, a primary bottleneck in scalable tool invocation is the precise selection of candidate tools from expansive libraries.
+Presenting thousands of tool descriptions to an LLM is often infeasible due to limited context windows, semantic noise from irrelevant options, and prohibitive computational latency.
+Current solutions typically rely on preliminary filtering via keyword or vector search; however, these methods frequently overlook query nuances and the inherent server-tool hierarchy within libraries (Fei et al., 2025).
+Treating the library as a flat list often leads to suboptimal filtering, where relevant tools are discarded while noise is retained, ultimately degrading selection accuracy.
+To address these limitations, we propose the Hierarchical Gaussian Mixture Framework (HGMF), a probabilistic pruning framework designed for scalable and structured tool selection.
+HGMF maps the query, servers, and tools into a unified semantic space and implements a hierarchical filtering process using Gaussian Mixture Models (GMMs).
+By first identifying relevant server clusters based on query likelihood and subsequently pruning tools within those clusters, HGMF yields a refined candidate set for final reranking by an LLM.
+Figure 1.
+Overall Pipeline.
+The process involves embedding the user query, servers, and tools into a unified semantic space.
+HGMF then performs a two-level pruning strategy using clustering at the server level to obtain a relevant server list, followed by clustering at the tool level within the selected servers to yield a refined list of candidate toolsets.
+Our main contributions are: (1) We propose HGMF, a novel hierarchical framework that leverages the server-tool architecture to decompose the complex search space for tool invocation.
+(2) We adapt GMMs for probabilistic soft clustering, which more effectively captures the semantic correspondence between queries and tool clusters compared to deterministic filters.
+(3) We introduce a GMM regularization technique, incorporating constraints such as inter-class separation and intra-class compactness, to enhance robustness and accuracy in sparse-data scenarios.
+(4) We demonstrate through rigorous evaluation that HGMF achieves state-of-the-art performance on the MCP-tools dataset, outperforming baselines by up to 40% in accuracy across eight LLMs while significantly reducing inference latency.
+1.1. Related Works
+1.1.1. Tool Augmentation in Large Language Models
+Integrating external tools into LLMs has emerged as a core paradigm for transcending inherent knowledge boundaries and empowering LLMs to tackle complex real-world tasks (Chen et al., 2024).
+From early retrieval-augmented generation (RAG) (Lewis et al., 2020) to the pioneering self-supervised tool invocation in Toolformer (Schick et al., 2023), recent studies have focused on deepening LLMs’ tool utilization capabilities, such as enhancing awareness of tool preconditions (Yang et al., 2024) or improving decision reliability through meta-verification and reflection learning (Chen et al., 2024).
+Comprehensive surveys have also appeared, systematically reviewing methods and challenges in tool learning for LLM-based agents (Xu et al., 2025a).
+1.1.2. Challenges and Existing Methods in Large-Scale Tool Selection
+Consequently, as tool libraries scale to thousands of APIs, efficient selection from vast repositories has become a central bottleneck.
+Standardized interfaces like the Model Context Protocol (MCP) (Anthropic, 2024) have facilitated unified interactions and ecosystem growth, yet this proliferation exacerbates the challenge by overwhelming traditional in-context learning approaches, rendering them infeasible in large-scale settings.
+To address this, the paradigm has shifted toward two-stage retrieval-ranking” strategies, such as random sampling in early MCP protocols (Fei et al., 2025), which yield low hit rates and prove impractical.
+Vector-based semantic search serves as a baseline for efficiency, but its flat” nature introduces noise from lexical overlaps in semantically similar tools, leading to errors (Fei et al., 2025).
+Advanced techniques, including Top-P sampling (Holtzman et al., 2019) and dynamic retrieval (Lumer et al., 2025), offer refinements, yet they largely neglect the inherent hierarchical structures within tool libraries.
+However, these applications primarily focus on flat” clustering of unstructured text.
+2. Methodology
+We propose the Hierarchical Gaussian Mixture Framework (HGMF), a three-stage process shown in Figure 1, to efficiently prune large toolsets for LLMs while mitigating context and noise issues.
+HGMF first creates semantic embeddings, then applies regularized, hierarchical GMM-based pruning, and finally uses an LLM for reranking.
+2.1. Semantic Embedding and Preprocessing
 The initial stage aims to represent all textual information—the user query, server descriptions, and tool descriptions—within a unified high-dimensional semantic space.
@@
 To ensure that the subsequent similarity and probability calculations are not biased by vector magnitudes, we apply L2 normalization to all generated embeddings.
-For any vector $\mathbf{v}$, its normalized counterpart $\hat{\mathbf{v}}$ is computed as:
-$\hat{\mathbf{v}}=\frac{\mathbf{v}}{\|\mathbf{v}\|_{2}}$ (1)
+For any vector $\mathbf{v}$, its normalized counterpart $\hat{\mathbf{v}}$ is computed as: $\hat{\mathbf{v}}=\frac{\mathbf{v}}{\|\mathbf{v}\|_{2}}$.
 This normalization projects all vectors onto the surface of a unit hypersphere, making cosine similarity equivalent to the dot product and stabilizing the clustering process.
-Fig.
-1: Overall Pipeline.
-The process involves embedding the user query, servers, and tools into a unified semantic space.
-HGMF then performs a two-level pruning strategy using clustering at the server level to obtain a relevant server list, followed by clustering at the tool level within the selected servers to yield a refined list of candidate toolsets.
-3.2 Hierarchical GMM-based Candidate Pruning
-This stage is the cornerstone of our method, designed to drastically reduce the search space by exploiting the inherent server-tool hierarchy.
-We employ a two-level pruning strategy using Gaussian Mixture Models (GMMs).
+2.2. Hierarchical GMM-based Candidate Pruning
+We employ a two-level strategy using Gaussian Mixture Models (GMMs) enhanced with inter-class separation and intra-class compactness regularization.
+These constraints are integrated into the GMM’s iterative fitting process to ensure robust and well-defined probabilistic clustering, especially on sparse data.
+Inter-class and Intra-class Regularization
+The regularization incorporates four key parameters: inter-cluster regularization ($\lambda_{\text{inter}}$), intra-cluster regularization ($\beta_{\text{intra}}$), balance coefficient ($w_{\text{balance}}$), and covariance regularization ($\text{reg}_{\text{covar}}$).
+The inter-class term enforces separation by penalizing proximity between cluster centers:
+$\mathcal{L}_{\text{inter}}=\lambda_{\text{inter}}\sum_{i=1}^{K}\sum_{j\neq i}\|\mu_{i}-\mu_{j}\|^{2}$
+where $\mu_{i}$ and $\mu_{j}$ represent the mean vectors of the $i$-th and $j$-th cluster centers, respectively.
+In the case of small sample sizes, cluster centers tend to aggregate, leading to blurred cluster boundaries.
+Inter-class separation ensures sufficient distinction between different classes.
+The intra-class regularization term is defined as:
+$\mathcal{L}_{\text{intra}}=\beta_{\text{intra}}\sum_{i=1}^{K}\text{tr}(\Sigma_{i})$
+where $\Sigma_{i}$ denotes the covariance matrix of the $i$-th cluster, and $\text{tr}(\cdot)$ represents the trace of a matrix.
+This term maintains reasonable cluster shapes, preventing excessive elongation or deformation.
+The balance term is defined as:
+$\mathcal{L}_{\text{balance}}=w_{\text{balance}}\cdot(\mathcal{L}_{\text{inter}}+\mathcal{L}_{\text{intra}})$
+This term balances the conflicting objectives of inter-class separation and intra-class compactness, preventing any single regularization term from dominating.
+The regularized covariance matrix is defined as:
+$\Sigma_{i}^{\text{reg}}=\Sigma_{i}+\text{reg}_{\text{covar}}\cdot I$
+where $I$ is the identity matrix.
+This regularization prevents the covariance matrix from becoming singular.
+Our method incorporates regularization constraints during the M steps of Gaussian model training, followed by iterative updates until convergence.
+We also set convergence thresholds and maximum iteration limits to control training accuracy and duration.
+Table 1.
+Average accuracy (%) of all methods across different LLMs.
+The best result in each row is in bold.
+Abbreviations: Tokeniz.
+= Tokenization-Based, CW.
+= Cluster-Weighted, Density = Density-Based, Hermes.
+= Openhermes.
+LLM MCP-zero HGMF Tokeniz. K-means CW. Density
+Mistral (7b) 80.95 91.32 58.65 82.75 73.21 78.35
+Gemma (7b) 79.52 79.19 38.43 62.05 56.04 56.77
+Llama3 (8b) 82.74 84.83 67.50 83.52 74.73 83.15
+Qwen (14b) 86.26 87.21 79.88 87.51 84.25 78.64
+Solar (10.7b) 80.36 87.99 69.61 75.85 78.28 75.56
+Hermes. (7b) 75.79 82.57 47.48 74.15 71.87 60.87
+Phi3 (3.8b) 59.48 62.58 40.92 54.80 55.74 55.86
+Phi4 (14b) 79.70 82.99 80.18 84.95 82.93 79.60
+Average 78.10 82.34 60.32 75.70 72.13 71.10
 Server-Level Pruning.
@@
 The likelihood score for the $k$-th cluster is given by its probability density function:
-$\mathcal{L}(k|\hat{\mathbf{v}}_{q})=\mathcal{N}(\hat{\mathbf{v}}_{q}|\bm{\mu}_{k},\bm{\Sigma}_{k})$ (2)
+(1) $\mathcal{L}(k|\hat{\mathbf{v}}_{q})=\mathcal{N}(\hat{\mathbf{v}}_{q}|\bm{\mu}_{k},\bm{\Sigma}_{k})$
 We rank all server clusters based on these likelihood scores in descending order and select the top-$N_{s}$ clusters.
@@
 This hierarchical process ensures that only tools associated with relevant servers are considered, leading to a highly refined and contextually coherent candidate pool.
-Fig.
-2: Performance comparison of our proposed method (HGMF) against five baselines across eight different Large Language Models (LLMs).
-Each subplot corresponds to a specific LLM, showing accuracy (Y-axis) as a function of sample size (X-axis, log scale).
-The results indicate that HGMF achieves state-of-the-art performance under most models and shot counts, with a particularly significant advantage at larger sample sizes.
-3.3 LLM-based Reranking and Final Selection
+Figure 2.
+Performance comparison of HGMF against five baselines across eight LLMs.
+Each subplot shows accuracy as a function of sample size (log scale).
+HGMF achieves state-of-the-art results in most scenarios, showing a clear advantage at larger sample sizes.
+2.3. LLM-based Reranking and Final Selection
 After pruning, we have a small, structured candidate set $\{(s_{i},t_{j})|s_{i}\in\mathcal{S}^{\prime},t_{j}\in\mathcal{T}^{\prime}\}$.
@@
 We then perform the final matching by computing the cosine similarity between the LLM-generated embeddings and the embeddings of all candidates in the pruned set.
-The server score for a candidate server $s_{i}\in\mathcal{S}^{\prime}$ is:
-$\text{score}(s_{i})=\cos(\mathbf{v}_{s}^{*},\hat{\mathbf{v}}_{s_{i}})$ (3)
-Similarly, the tool similarity for a candidate tool $t_{j}\in\mathcal{T}^{\prime}$ is:
-$\text{sim}(t_{j})=\cos(\mathbf{v}_{t}^{*},\hat{\mathbf{v}}_{t_{j}})$ (4)
+The server score for a candidate server $s_{i}\in\mathcal{S}^{\prime}$ is: $\text{score}(s_{i})=\cos(\mathbf{v}_{s}^{*},\hat{\mathbf{v}}_{s_{i}})$.
+Similarly, the tool similarity for a candidate tool $t_{j}\in\mathcal{T}^{\prime}$ is: $\text{sim}(t_{j})=\cos(\mathbf{v}_{t}^{*},\hat{\mathbf{v}}_{t_{j}})$.
 To determine the final ranking, we compute a combined score for each server-tool pair $(s_{i},t_{j})$ that accounts for both server-level and tool-level matching quality.
 The final score is calculated as:
-$\begin{split}\text{FinalScore}(s_{i},t_{j})={}&(\text{score}(s_{i})\times\text{sim}(t_{j}))\\ &\times\max(\text{score}(s_{i}),\text{sim}(t_{j}))\end{split}$ (5)
+(2) $\begin{split}\text{FinalScore}(s_{i},t_{j})={}&(\text{score}(s_{i})\times\text{sim}(t_{j}))\\ &\times\max(\text{score}(s_{i}),\text{sim}(t_{j}))\end{split}$
 This formula gives higher weight to pairs where both the server and tool are a strong match, and it particularly rewards candidates where at least one component has an exceptionally high similarity score.
 The server-tool pair with the highest final score is selected as the output of our HGMF framework.
-4 Experiments
-4.1 Experiment Configuration
-Dataset
-The dataset employed in this study is MCP-tools, introduced by Fei et al.
-[13].
-It comprises 2,797 tools collected from 308 servers within the official MCP repository.
-For our semantic matching task, we preprocessed all textual content by generating vector embeddings using the all-MiniLM-L6-v2 sentence-transformer model [14].
-The original data, along with these pre-computed embeddings, were consolidated and stored in a single JSON file.
-Table 1: Average accuracy (%) of all methods across different LLMs. The best result in each column is in bold. Full results are in the Appendix.
-Method Mistral (7b) Gemma (7b) Llama3 (8b) Qwen (14b) Solar (10.7b) Openhermes (7b) Phi3 (3.8b) Phi4 (14b) Average
-MCP-zero 80.95 79.52 82.74 86.26 80.36 75.79 59.48 79.70 78.10
-HGMF (Ours) 91.32 79.19 84.83 87.21 87.99 82.57 62.58 82.99 82.34
-Tokenization-Based 58.65 38.43 67.50 79.88 69.61 47.48 40.92 80.18 60.32
-K-means 82.75 62.05 83.52 87.51 75.85 74.15 54.80 84.95 75.70
-Cluster-Weighted 73.21 56.04 74.73 84.25 78.28 71.87 55.74 82.93 72.13
-Density-Based 78.35 56.77 83.15 78.64 75.56 60.87 55.86 79.60 71.10
-Implementation Details
-Our implementation uses Python 3.10 and PyTorch 2.7.1 (CUDA 12.6), executed on NVIDIA RTX 4090 GPUs.
-Key hyperparameters are: $k=4$ for top-k retrieval, a maximum of 10 tools indexed per server, and random sampling for sample sizes under 10 to mitigate noise.
-We evaluate on 10 exponentially scaled sample sizes ranging from 1 to 2797.
-Evaluation Metrics
-We evaluate model performance using Exact Match Accuracy. A prediction is considered correct only if both the retrieved server and tool perfectly match the ground truth. The metric is calculated as:
-$\text{Accuracy}=\frac{\text{\# Correct Predictions}}{\text{\# Total Test Cases}}$
-We report accuracy for the entire test set (Overall) and for each individual sample size tier ($s$).
-Baselines.
-We compare our method against the following baselines:
-  • MCP-zero [13]: The current state-of-the-art, which uses random sampling without replacement.
-  • Tokenization-Based: A naive approach that samples based on tokenizing the input.
-  • K-means: A standard clustering approach where we sample from the $k$ clusters closest to the query embedding.
-  • Cluster-Weighted: An extension of K-means that assigns sampling weights to tools within selected clusters based on their cosine similarity to the query.
-  • Density-Based Clustering: A method that groups servers and tools based on their intrinsic embedding similarities without a predefined number of clusters.
-Fig. 3: Accuracy gain of the HGMF method over the MCP-zero baseline. The results show that HGMF’s performance advantage steadily increases with the sample size, demonstrating its superior scalability.
-4.2 Result Analysis
-Table 1 and Fig.
-2 present a comprehensive comparison of our proposed method, HGMF, against five baseline methods across eight different LLMs and four sample sizes.
+3. Experiments
+3.1. Experiment Configuration
+Figure 3.
+Accuracy comparison of HGMF with and without regularization.
+The regularized model achieves significant gains (14-28%) in low-shot scenarios by mitigating cluster instability, leading to more stable and accurate performance across all sample sizes.
+[…]
+Figure 4.
+Accuracy gain of the HGMF method over the MCP-zero baseline.
+The results show that HGMF’s performance advantage steadily increases with the sample size, demonstrating its superior scalability.
+3.2. Result Analysis
+Overall Performance
+Figure 2 present a comprehensive comparison of our proposed method, HGMF, against five baseline methods across eight different LLMs and four sample sizes.
 The results demonstrate that HGMF consistently achieves superior accuracy, particularly in high-shot settings.
-Fig.
-3 shows the accuracy improvement of our proposed HGMF method over the MCP-zero baseline across eight LLMs and four sample sizes.
+Figure 4 shows the accuracy improvement of our proposed HGMF method over the MCP-zero baseline.
 Positive values indicate performance gains, while negative values indicate underperformance.
-Key observations are summarized as follows:
-Overall Performance.
-HGMF attains the highest average accuracy (82.34%) across all models, outperforming the strong baseline MCP-zero (78.10%) by over 4 percentage points.
-It achieves the best results on 6 out of 8 models, including Mistral:7b, Llama3:8b, Qwen:14b, Solar:10.7b, Openhermes:7b, and Phi3:3.8b.
+In summary, HGMF attains the highest average accuracy (82.34%) across all models, outperforming the strong baseline MCP-zero (78.10%) by over 4 percentage points.
+It achieves the best results on 6 out of 8 models.
 This confirms the effectiveness of our hierarchical clustering and pruning strategy in selecting high-quality tool candidates.
-Comparison with Clustering-Based Baselines:
-While K-means and Density-Based Clustering exhibit moderate improvements over Tokenization-Based sampling, they fall short of HGMF.
-Notably, Cluster-Weighted performs inconsistently, likely due to its reliance on local similarity scores without hierarchical context.
-In contrast, HGMF integrates both server-level and tool-level clustering, enabling more robust and context-aware selection.
-Scalability with Sample Size:
-HGMF’s performance gain increases consistently with more samples—from 41-shot (light gray) to 2797-shot (dark blue)—demonstrating strong scalability and its ability to extract high-quality candidates from larger tool pools.
-Low-Shot Limitations:
-In 41-shot settings, HGMF may underperform due to insufficient context for effective clustering. Notably, it lags behind MCP-zero on models like Gemma:7b and phi4:14b, with up to 30% accuracy drop.
-High-Shot Superiority:
-At larger sample sizes ($\geq$401-shot), HGMF consistently outperforms the baseline.
-At 2797-shot, it delivers significant gains—over 40 percentage points on Openhermes:7b and 30+ points on phi4:14b, highlighting its strength in distilling relevant tools from noisy candidates.
-Model Sensitivity:
-Response to HGMF varies by model. Gemma:7b shows large fluctuation, while models like Mistral:7b and Qwen:14b exhibit stable improvements across all settings.
-5 Conclusion
-To address the challenges of context limitations and noise interference when Large Language Models select tools from large, hierarchical libraries, this paper introduces the Hierarchical Gaussian Mixture Framework (HGMF).
-The framework leverages the library’s inherent ”server-tool” structure, performing efficient probabilistic pruning on the vast toolset through two-stage Gaussian Mixture Model (GMM) clustering to produce a compact, high-relevance candidate set for the LLM.
-Experiments confirm HGMF’s effectiveness and scalability.
-It achieves higher accuracy than baseline methods like random sampling across multiple LLMs.
-Critically, its performance advantage grows with the size of the tool library, demonstrating that the framework can effectively handle the challenges of large-scale, real-world toolsets.
-Through this intelligent filtering, HGMF not only improves selection accuracy but also reduces the LLM’s inference burden.
-The core contribution of this work is an efficient and viable paradigm for large-scale tool invocation.
-By integrating hierarchical information into the pruning process, HGMF successfully decomposes a complex selection problem into more manageable sub-tasks.
-Although its performance is limited with very small toolsets, this work points toward future research, such as exploring adaptive clustering parameters and extending the framework to more complex, graph-structured tool libraries.
+High-Shot Superiority
+As shown in Figure4, at larger sample sizes ,HGMF consistently outperforms the baseline.
+It delivers significant gains—over 40 percentage points on Openhermes:7b and 30+ points on phi4:14b.
+highlighting its strength in distilling relevant tools.
+Effect of Regularization
+An ablation study (Figure 3) confirms that our proposed regularization technique effectively mitigates performance instability on sparse data.
+While the baseline model without regularization struggles at low sample sizes (57-61% accuracy), the fully regularized HGMF yields an average accuracy gain of 21.65 percentage points in these low-shot regimes by preventing unstable clustering.
+4. Conclusion
+To tackle context limitations and noise in Large Language Models (LLMs) selecting tools from large, hierarchical libraries, this paper introduces the Hierarchical Gaussian Mixture Framework (HGMF).
+It exploits the library’s ”server-tool” structure, using two-stage Gaussian Mixture Model (GMM) clustering for efficient probabilistic pruning to generate a compact, highly relevant candidate set for the LLM.
+Experiments validate HGMF’s effectiveness and scalability, achieving higher accuracy than baselines like random sampling across various LLMs.
+Its advantage increases with tool library size, proving suitability for large-scale, real-world toolsets.
+By smart filtering, HGMF enhances selection accuracy and reduces LLM inference load.
+The core contribution is an efficient paradigm for large-scale tool invocation, integrating hierarchical information to break down complex selection into manageable sub-tasks.
+Though limited on very small toolsets, it suggests future directions like adaptive clustering parameters and extensions to graph-structured libraries.